Determine the volume of the torgus generated by

the revolving circular region enclosed by

(x − 3)²+ y² = 1 about the y-axis.


1
Expert's answer
2021-06-07T18:36:52-0400

Given circle is (x-3)2 + y2 = 1

Centre of circle is (3,0) and radius is 1.

We know that Volume, V = "\pi" "\int"ab (R2 - r2) dy

Here a = -1 and b = 1

R = 3+(1-y2)(1/2) and r = 3-(1-y2)(1/2)

Now V ="\Pi" "\int"-11 [9+(1-y2)+6(1-y2)(1/2)]-[9+(1-y2)-6(1-y2)(1/2)]

V = 12 "\Pi" "\int"-11 (1-y2)(1/2) dy

We know that"\int"(a2-x2)dx=[(1/2)x(a2-x2)(1/2)+(1/2)a2 sin-1(x/a) + C

So, V =12"\Pi" [(1/2)(1-(1)2)(1/2)+(1/2)sin-1(1)-(1/2)(1-(-1)2)(1/2)-(1/2)sin-1(-1)]

V = 12"\Pi" * "\Pi" = 12 "\Pi" 2 Units3